OpenStudy (anonymous):

Factor the trinomial: 2x2 - 5x - 3 (2x + 1)(2x - 3) (2x + 1)(x - 3) (2x - 1)(x - 3) (2x - 1)(2x + 3)

OpenStudy (anonymous):

I think its B) 2x * x = 2x^2 2x * -3 = -6x 1*x = x 1*-3 = -3 combine: 2x^2 -6x + x - 3 2x^2 -5x - 3

OpenStudy (anonymous):

Thanks so much

OpenStudy (anonymous):

The above work is just checking to make sure it equals the original trinomial:)

OpenStudy (anonymous):

your welcome:)

OpenStudy (anonymous):

@jazy and @TuringTest for helping me